The current one-year Treasury rate is 4.07% and the two-year rate is 5.07%. Assume expectations theory. What is the one-year rate expected one year from today (E(1r2)). Enter in percent form without the percent sign.
Answer #1

One year rate expected one year from today =(1+rate year 2)^(2)/(1+rate 1)-1 =(1+5.07%)^2/(1+4.07%)-1 =6.08%

one-year rate expected one year from today (E(1r2) =6.08%